OpenStudy (anonymous):

A giant pie is created in an attempt to break a world record for baking. The pie is shown below:

OpenStudy (anonymous):

What is the area of the slice of pie that was cut, rounded to the nearest hundredth

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

The area of the whole circle is \[ (7.5)^2\pi \]

OpenStudy (anonymous):

The slice area is \[ \frac{(7.5)^2\pi}{360} \times 45 \]

OpenStudy (anonymous):

It is about 22.0893

OpenStudy (anonymous):

wow! thanks for not just giving me the answer and helping me work thru it

OpenStudy (anonymous):

YW